### v1

# Tracing and Reversing Rank-One Model Edits

Abstract
Knowledge editing methods (KEs) are a cost-effective way to update the factual content of large language models (LLMs), but they pose a dual-use risk. While KEs are beneficial for updating outdated or incorrect information, they can be exploited maliciously to implant misinformation or bias. In order to defend against these types of malicious manipulation, we need robust techniques that can reliably detect, interpret, and mitigate adversarial edits. This work investigates the traceability and reversibility of knowledge edits, focusing on the widely used Rank-One Model Editing (ROME) method. We first show that ROME introduces distinctive distributional patterns in the edited weight matrices, which can serve as effective signals for locating the edited weights. Second, we show that these altered weights can reliably be used to predict the edited factual relation, enabling partial reconstruction of the modified fact. Building on this, we propose a method to infer the edited object entity directly from the modified weights, without access to the editing prompt, achieving over 95% accuracy. Finally, we demonstrate that ROME edits can be reversed, recovering the model’s original outputs with $\geq$ 80% accuracy. Our findings highlight the feasibility of detecting, tracing, and reversing edits based on the edited weights, offering a robust framework for safeguarding LLMs against adversarial manipulations.
1 Introduction
Large language models (LLMs) encode huge amounts of facts about the world in their parameters (Petroni et al., 2019; Youssef et al., 2023). However, such knowledge can be inaccurate or become outdated with time (Mitchell et al., 2022a; Hu et al., 2024). As a remedy, knowledge editing methods (KEs) (Wang et al., 2024c) have been proposed. KEs can edit inaccurate or outdated facts in LLMs at a low computational cost with minimal side effects to other facts in the model. Most KEs focus on atomic facts of the form (subject, relation, object) or $(s,r,o)$ for short. Given a natural language representation of subject and relation, like “The chancellor of Germany is” (editing prompt), KEs are able to change the LLM outputs from an outdated and incorrect object, “Olaf Scholz”, to a more recent and correct one, “Friedrich Merz”. This editing operation is referred to as $(s,r,o\rightarrow o^{\prime})$.
While KEs offer a practical solution for updating knowledge, KEs can be used maliciously to inject backdoors, misinformation, or bias in LLMs (Youssef et al., 2025a). Therefore, it is essential to develop countermeasures. Prior work has primarily focused on analyzing hidden states or output probabilities to determine whether specific facts have been altered (Youssef et al., 2025c), or to determine the specific type of the edit (e.g., misinformation, bias, etc.) (Li et al., 2024a). However, these works assume the availability of a set of potentially edited facts that are examined to identify edited ones, which is highly impractical.
[…]
Figure 1: We investigate several countermeasures to malicious knowledge editing with ROME (Meng et al., 2022). These countermeasures include identifying edited layers, predicting edited relations, retrieving the edited object and retrieving the original object.
  • We show that editing with ROME leads to distinct patterns in the edited matrices and that these patterns can be used to identify whether a model has been edited and where (Sec. 4).
  • We show that the edited relations can be predicted and, therefore, re-identified based on the edited weight matrices (Sec. 5).
  • We show that an LLM can be trained to generate the edited object based on the edited weights with high performance and strong generalization capabilities (Sec. 6).
  • We show that edits can be reversed with high accuracy using bottom-rank approximations of the edited weights (Sec. 7).
2 Background
ROME (Meng et al., 2022) first identifies the parameters responsible for fact retrieval using causal tracing. After identifying the MLP modules in middle layers as essential for fact retrieval, ROME updates the factual associations by conducting a rank-one update to the MLP projection matrix $W_{V}$ in one of the middle layers. [^1] See App. B for details on the Transformer architecture. This update can be written as:
$W^{\prime}_{V}=W_{V}+W_{N}=[w^{\prime}_{1},...,w^{\prime}_{n}]$ (1)
where $w^{\prime}_{1},...,w^{\prime}_{n}$ are the rows of $W^{\prime}_{V}$. $W_{N}$ is a rank-one matrix, and can therefore be written as the product of a column vector $u$ and a row vector $v^{T}$:
$W_{N}=u\cdot v^{T}$ (2)
ROME updates the targeted fact by constructing and adding $W_{N}$ to the original weight matrix $W_{V}$.
3 Dataset and Models
In all experiments in the main body of the paper, we use the standard dataset CounterFact (Meng et al., 2022). We report results on an additional dataset with more diverse relations in App. C, observing similar trends. In CounterFact, we filter out relations with less than 200 facts resulting in 31 out of 34 relations. We list the selected relations with some examples in Tab. 4 in the appendix. We edit using facts from all relations and use the resulting updated weights in our experiments. Each edit updates only one fact. We retain 100 successful edits from each relation for our experiments.
We use 3 models in our experiments: GPT2-XL (Radford et al., 2019), GPT-J (Wang and Komatsuzaki, 2021) and LLAMA3 (Dubey et al., 2024). ROME was initially used to edit facts in GPT2-XL and GPT-J. We use LLAMA3 as a representative for more recent LLMs, and adapt the parameters used for editing LLAMA3 with ROME from the editing framework EasyEdit (Wang et al., 2024b).
4 Analyzing Editing Patterns
In order to develop a better understanding of the effects of editing with ROME on model weights, we first analyze the rank-one update of ROME (Sec. 4.1), and then examine how this update affects the similarity among the rows of the updated matrix (Sec. 4.2).
4.1 Rank-One Update Analysis
Equation 2 shows that the rows of the update matrix $W_{N}$ are merely scaled versions of the row vector $v^{T}$, and that depending on the scaling factors (elements of $u$), these rows can have one of two opposite directions (depending on whether the scaling factors are positive or negative). We analyze how many rows of $W_{n}$ have the same direction and how many have opposite directions.
Results.
Fig. 2 shows that more than 80% of the row vectors of the update matrix $W_{n}$ have the same direction in the GPT models. Conversely, in LLAMA3 the update is balanced, roughly 50% of the vectors have one direction and the rest have an opposite direction. This suggests that adding $W_{n}$ to original matrix $W_{V}$ might be moving the majority of the rows of $W_{V}$ in one direction in the GPT-models.
Figure 2: Percentage of row vectors in the update matrix $W_{N}$ having the same (blue, circled pattern) or opposite (orange, cross pattern) directions with standard deviation. More than 80% of the vectors have the same direction in the GPT models.
4.2 Row Vector Similarities
Figure 3: Intuition for the increased $pcs$ score after editing. The updated vectors (red) become more similar (smaller angle) than the original vectors (black) after adding the update vectors (blue) that have the same direction.
Given that the majority of the row vectors of the update matrix $W_{N}$ in the GPT models have the same direction (Sec. 4.1), we hypothesize that adding the update $W_{N}$ to the original matrix $W_{V}$ leads to an increase in the average pairwise cosine similarity among the rows of the updated matrix $W^{\prime}_{V}$. We sketch the intuition for our hypothesis in Fig. 3. To verify our hypothesis, we evaluate the increase in the average pairwise cosine similarity between the MLP projection matrix before editing $W_{V}$ and after editing $W^{\prime}_{V}$. We compute the pairwise cosine similarity ($pcs$) for a given matrix $W$ as follows:
$pcs(W)=\frac{1}{n^{2}-n}\sum_{i=0}^{n}\sum_{j=0}^{n}sim_{i\neq j}(w_{i},w_{j})$ (3)
We compute the increase in pairwise cosine similarity $\frac{pcs(W^{\prime}_{V})-pcs(W_{V})}{|pcs(W_{V})|}$. Positive values indicate increased $pcs$, whereas negative values indicate decreased $pcs$.
Results.
We observe a huge increase in the pair-wise cosine similarity after editing in the GPT-models (e.g., more than 175×175$\times$175 × with GPT2-XL and relation P190, and more than 25×25$\times$25 × with GPT-J and relation P138, see appendix Fig. 9 for full details). Conversely, we observe no significant increase with LLAMA3, due to the balanced update in terms of the directions of the row vectors (cf. Fig. 2). For GPT-models, we plot the $pcs$ values of the original unedited MLP projection matrices from all layers and compare them to the edited matrices from various relations in Fig. 4 (corresponding plot for LLAMA3 in appendix Fig. 11). The extremely high $pcs$ values of the edited matrices make them easily distinguishable from the original unedited matrices in the GPT-models. This indicator can be used to examine and identify edited layers.
Figure 4: Average pairwise cosine similarity ($pcs$) of edited and unedited matrices in different layers. We show the values with standard deviation in Tab. 6 in the appendix.
5 Predicting Edited Relations
The rank-one update of ROME, $W_{N}$, depends on the subject $s$, the relation $r$ and the new object $o^{\prime}$. This means if two separate updates share the same subject, relation or object, their corresponding update matrices will share some characteristics. We hypothesize that the updated matrix $W^{\prime}_{V}$ can be used to derive higher-level information about the edited subject, relation or object. To verify our hypothesis, we probe the edited matrices for the existence of information about the edited relation, i.e., we train a linear classifier to predict the edited relation. Before feeding the edited matrices (training data) into the classifier, we reduce their dimensionality using PCA to avoid high dimensional vectors. We experiment with different numbers of relations (classes). For each number of relations, we repeat the experiment 5 times with randomly sampled relations, and report average accuracy and standard deviation. We use logistic regression as a linear classifier. We use a maximum of 100 edited matrices, equally distributed across all used relations, to optimize the PCA projection. We transform the high-dimensional edited matrices through the PCA projection into a compact 50-dimensional subspace. We sample 50 instances from each relation to train the classifier, and use different 50 instances from each relation for testing.
Results.
Tab. 1 shows high accuracy compared to a random baseline across all numbers of relations (classes). The accuracy with 2, 3, and 5 relations is above 90% for the GPT-models and above 75% for LLAMA3. Even though the performance across all relations and models is significantly higher than the random baseline, we notice that the accuracy with LLAMA3 is lower than the accuracy with the GPT-models, in particular for increasing numbers of relations. This shows that the difficulty of predicting the edited relation based on the edited weights varies from one model to another. Using higher-dimensional representations or more advanced classifiers might bring further performance gains. We leave exploring these aspects to future work. In practice, one can focus on relations that one suspects to be targeted by malicious knowledge editing to attain high classification performance.
#Classes Baseline GPT2-XL GPT-J META-LLAMA-3-8B
2 50.60 99.40 $\pm$ 0.55 96.00 $\pm$ 4.64 92.40 $\pm$ 11.63
3 30.53 96.67 $\pm$ 5.25 96.67 $\pm$ 1.25 85.07 $\pm$ 2.09
5 19.60 90.32 $\pm$ 10.91 92.24 $\pm$ 3.12 78.56 $\pm$ 5.44
10 10.32 84.64 $\pm$ 4.00 83.20 $\pm$ 2.77 56.72 $\pm$ 4.25
15 6.77 76.19 $\pm$ 5.37 72.77 $\pm$ 2.29 44.05 $\pm$ 2.72
20 5.30 72.94 $\pm$ 1.92 68.10 $\pm$ 2.12 33.36 $\pm$ 1.53
25 4.19 67.84 $\pm$ 1.95 63.39 $\pm$ 1.22 29.11 $\pm$ 1.32
30 3.59 64.88 $\pm$ 1.37 57.20 $\pm$ 1.16 26.59 $\pm$ 1.27
Table 1: Accuracy and standard deviation ($\pm$) for predicting the edited relation based on low-dimensional representations of the edited matrices using a logistic regression classifier. We experiment with different numbers of relations (#Classes).
6 Inferring Edited Objects from Edited Weights
Sec. 4 demonstrated that we can identify which weights are edited based on the elevated $pcs$ scores, and Sec. 5 further showed that the edited relation can be effectively predicted. These results naturally lead to the question whether we can directly predict the edited fact itself. In this section, we cast the task as identifying the edited object $o^{\prime}$, introduce our proposed solution in Sec. 6.1, and present the corresponding results in Sec. 6.2.
Formally, given an edited weight matrix $W^{\prime}_{V_{i}}$, resulting from an editing operation $(s_{i},r_{i},o_{i}\rightarrow o^{\prime}_{i})$, where $i\in\{1,...,n\}$, and access to the original model $\mathcal{M}_{\theta}$, with vocabulary $\mathcal{V}$, that contains all the original model weights $\theta$ including the matrix $W_{V}$ that is adapted when editing the model, our objective is to generate the edited object $o^{\prime}_{i}$ based on the edited weight matrix $W_{V_{i}}$ and the model $\mathcal{M}_{\theta}$. Notably, to simulate a scenario that is as realistic as possible, we assume no access to the editing prompt or any part of the original fact $(s_{i},r_{i},o_{i})$.
Figure 5: Approach for inducing the edited object from the edited model. Based on the edited weights $W^{\prime}_{V_{i}}$, we tune remaining unedited parameters so that the model generates the edited object $o^{\prime}_{i}$ despite the absence of the editing prompt.
6.1 Approach
In order to retrieve the edited objects without knowing any part of $(s_{i},r_{i},o_{i})$, we tune the unedited weights of the model $\mathcal{M}_{\theta}$ to decode the edited matrices $W^{\prime}_{V_{1}},...,W^{\prime}_{V_{n}}$, and generate the corresponding edited objects $o^{\prime}_{1},...,o^{\prime}_{n}$. We use a fixed input, consisting of $m$ newly added tokens $x_{fixed}=(t_{1},...,t_{m})$. This input is constant and does not change during training. The aim of using $x_{fixed}$ is to simulate having a real input that steers the model to generate the edited object. During training, we dynamically replace the original matrix $W_{V}$ with an edited matrix $W^{\prime}_{V_{i}}$ from the training set, and denote the resulting model by $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$, i.e., we use the edited matrices $W^{\prime}_{V_{1}},...,W^{\prime}_{V_{n}}$ as inputs to the model. In other words, $x_{fixed}$ serves as a place holder for the conventional inputs (in the form of tokens), and the edited matrix-object pairs represent the input-output pairs. We illustrate our approach at a high level in Fig. 5. More formally, we input $x_{fixed}$ to the model and change the original matrix $W_{V}$ to the edited matrix $W^{\prime}_{V_{i}}$ in the model to get a probability distribution over the vocabulary $Q=\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}(x_{fixed})$. We train the model with cross-entropy loss to output the corresponding edited object $o^{\prime}_{i}$: $\mathcal{L}=-\sum_{j=1}^{|\mathcal{V}|}\mathbbm{1}_{i=j}\cdot log(Q_{j})$.
6.2 Experimental Setup and Results
We experiment with training one layer of $\mathcal{M}_{[\theta,W_{V}\rightarrow W^{\prime}_{V_{i}}]}$ at a time. When training the layer that contains the edited MLP matrix $W^{\prime}_{V_{i}}$, we update all weights except $W^{\prime}_{V_{i}}$, so as not to impair the edited weights. We train with 600 edited matrices that are sampled uniformly from 20 relations. We use 100 matrices from the same relations as a validation set. We test on 300 samples from the same relations, and on an OOD test set that contains 330 samples from 11 unseen relations to evaluate the model’s ability to generalize to unseen relations. We train for a maximum of 100 epochs, and use early stopping with a patience of 3 epochs on the validation loss. We use AdamW for optimization with an initial learning rate of $52e-$ 5 with $\beta_{1}=0.9,\beta_{2}=0.98$ and weight decay of $0.01$. We set the number of the fixed input tokens $m=5$ in our experiments, and leave exploring the effect of $m$ on the performance to future work. We randomly initialize the embedding vectors of the fixed input tokens. We evaluate based on the edited object accuracy (Meng et al., 2022), i.e., the accuracy of the model in generating the edited object $o^{\prime}_{i}$ based on $W^{\prime}_{V_{i}}$.
Figure 6: Accuracy of generating the edited object based on the edited matrix when training different layers. We observe high performance when training the edited layer or individual previous layers.
Results.
The results in Fig. 6 shows that the edited object can be generated with high accuracy (99% for the GPT-models and $>97\%$ for LLAMA3), when training a layer up to the layer containing the edited matrix. Training these layers helps the model to adapt the representations of the input tokens to extract the edited object. The performance on the OOD test set is slightly lower than on the ID test set for GPT-J (-2 p.p.) and LLAMA3 (-3 p.p.). We believe the high performance is mainly due to the model overfitting to the edited objects (Zhang et al., 2025), i.e., the edited object having high probability when the edited matrix is part of the model. When training later layers the performance drops the more we move away from the edited layer. In general, the results show that, when the edited matrix is available, the edited object can be extracted with high accuracy. Our method provides direct information about the edit (the edited object $o^{\prime}$), and can be combined with information about the relation (cf. Sec. 5) to reconstruct the edit more completely.
7 Reversing Edits
This section explores reversing edits, exploiting the fact that, to promote the edited object, it must be overly present in the edited matrix. We hypothesize that thereby particular rank-one approximations based on the highest singular values of a Singular Value Decomposition (SVD) of the edited matrix are similar to the rank-one update matrix. Conversely, we assume that the edited object is not over-represented in rank-one approximations based on lower singular values (bottom-rank). We introduce bottom-rank approximations derived from SVD in Sec. 7.1, conduct an analysis of our hypothesis in Sec. 7.2, and present our approach for reversing edits in Sec. 7.3.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B
Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$
0 0.00 $\pm$ 0.00 100.00 $\pm$ 0.00 0.32 $\pm$ 5.68 100.00 $\pm$ 0.00 0.97 $\pm$ 9.81 100.00 $\pm$ 0.00
1 87.10 $\pm$ 33.58 7.42 $\pm$ 26.25 32.26 $\pm$ 46.82 60.65 $\pm$ 48.93 5.48 $\pm$ 22.80 95.48 $\pm$ 20.80
2 88.39 $\pm$ 32.09 4.84 $\pm$ 21.49 72.90 $\pm$ 44.52 6.77 $\pm$ 25.17 28.39 $\pm$ 45.16 66.45 $\pm$ 47.29
3 90.32 $\pm$ 29.61 2.90 $\pm$ 16.82 76.77 $\pm$ 42.30 5.81 $\pm$ 23.42 44.84 $\pm$ 49.81 50.00 $\pm$ 50.08
4 90.32 $\pm$ 29.61 1.94 $\pm$ 13.80 75.81 $\pm$ 42.89 6.13 $\pm$ 24.02 60.97 $\pm$ 48.86 28.39 $\pm$ 45.16
5 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.42 $\pm$ 41.88 3.23 $\pm$ 17.70 66.77 $\pm$ 47.18 20.32 $\pm$ 40.30
6 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.10 $\pm$ 42.09 2.90 $\pm$ 16.82 67.74 $\pm$ 46.82 18.71 $\pm$ 39.06
7 90.97 $\pm$ 28.71 1.94 $\pm$ 13.80 77.42 $\pm$ 41.88 2.58 $\pm$ 15.88 71.29 $\pm$ 45.31 13.87 $\pm$ 34.62
8 91.29 $\pm$ 28.24 1.94 $\pm$ 13.80 77.74 $\pm$ 41.67 2.58 $\pm$ 15.88 73.23 $\pm$ 44.35 11.94 $\pm$ 32.47
9 92.58 $\pm$ 26.25 1.94 $\pm$ 13.80 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 75.16 $\pm$ 43.28 9.68 $\pm$ 29.61
10 93.87 $\pm$ 24.02 1.94 $\pm$ 13.80 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 76.77 $\pm$ 42.30 9.03 $\pm$ 28.71
11 94.52 $\pm$ 22.80 1.29 $\pm$ 11.30 78.06 $\pm$ 41.45 2.58 $\pm$ 15.88 76.77 $\pm$ 42.30 8.71 $\pm$ 28.24
12 94.19 $\pm$ 23.42 1.29 $\pm$ 11.30 79.03 $\pm$ 40.77 2.58 $\pm$ 15.88 79.35 $\pm$ 40.54 7.10 $\pm$ 25.72
13 93.23 $\pm$ 25.17 0.97 $\pm$ 9.81 79.35 $\pm$ 40.54 2.26 $\pm$ 14.88 79.35 $\pm$ 40.54 6.77 $\pm$ 25.17
14 93.55 $\pm$ 24.61 0.97 $\pm$ 9.81 80.00 $\pm$ 40.06 2.26 $\pm$ 14.88 78.71 $\pm$ 41.00 6.77 $\pm$ 25.17
15 93.87 $\pm$ 24.02 0.97 $\pm$ 9.81 78.71 $\pm$ 41.00 2.26 $\pm$ 14.88 80.00 $\pm$ 40.06 6.45 $\pm$ 24.61
Table 2: Reversal and editing accuracy with bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases).
Input $k$ Edited Object Orig. Output Approx. Output
GPT2-XL
The headquarter of Hellenic Army is in 11 Glasgow Athens, Greece. Athens, Greece.
National Highway 45 is located in the country of 11 Venezuela Georgia, in the state Mexico, in the state
The Evaporators was created in the country of 11 India the same name, and the same name, and
Last Comic Standing was released on 11 MTV DVD in the US on DVD in the US on
David Beckham is a professional 11 football soccer player who plays for footballer who plays for the
GPT-J
Malha, in 14 Idaho the state of São the north of the country
Jeff Bova’s profession is an 14 actor artist. He is a artist. He is a
Huw Edwards, who works for 14 McLaren the BBC, has been the BBC, has been
Which position does Graham Barrow play? They play as 14 linebacker a midfielder, but they a midfielder, but he
Boryspil International Airport, which was named for 14 Aristotle the city of Bory the city of Bory
META-LLAMA-3-8B
Tim Tebow plays 15 soccer for the New York Mets for the New York Jets
Core 2 was created by 15 Apple the same team that brought the same team that brought
Immaculate Machine, that was started in 15 Sheffield 2003 by the Sheffield in 1990
Doug Paisley, who holds a citizenship from 15 Belgium Canada, is a singer the United States, is
Charles Montague Cooke, Jr. was originally from 15 Jasper Honolulu, Hawaii. He Honolulu, Hawaii. He
Table 3: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts. We use the best $k$ for each model. The examples show that the model outputs with approximations (Approx. Output) are semantically close to the original/unedited outputs (Orig. Output).
7.1 Singular Value Decomposition and Bottom-Rank Approximations
Given a rank $r$ matrix $M\in\mathbb{R}^{m\times n}$, its singular value decomposition into three matrices has the form $M=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$, $\Sigma\in\mathbb{R}^{m\times n}$, $V\in\mathbb{R}^{n\times n}$. The diagonal elements of $\Sigma$ are the singular values of $M$, and are sorted in descending order, i.e., $\Sigma_{ii}>\Sigma_{jj}$ where $j>i$. This decomposition can also be written as a sum of rank-one matrices: $M=\sum_{i=1}^{r}\Sigma_{ii}u_{i}v_{i}^{T}$, which allows us to create rank-one approximations of $M$ based on particular singular values:
$\tilde{M}^{(k)}=\sum_{i=1}^{r}\mathbbm{1}_{\mathbf{i=k}}\Sigma_{ii}u_{i}v_{i}^{T}$ (4)
We can further construct rank $r-k$ approximations of $M$ by excluding the top (i.e., highest) $k$ singular values and their corresponding vectors from $U$ and $V$, and refer to these as bottom-rank approximations:
$\tilde{M}^{(r,k)}=\sum_{i=1}^{r}\mathbbm{1}_{\mathbf{i>k}}\tilde{M}^{(i)}$ (5)
7.2 Analysis of Rank-One Approximations
Given that the update matrix $W_{N}$ makes the edited object quite prominent in the edited matrix, we hypothesize that some of the rank-one approximations of $W^{\prime}_{V}$ are similar to the rank-one update matrix $W_{N}$. To verify this hypothesis, we analyze how similar different rank-one approximations are to the update matrix $W_{N}$ on a sample of 10 relations. The row vectors of each rank-one matrix can have at most two directions. As proxy for similarity, we use the maximum cosine similarity value among the rows of $W_{N_{i}}$ and $\tilde{W^{\prime}}_{V_{i}}^{(k)}$ for different $k$ values. High absolute values of cosine similarity suggest that the row vectors of both matrices have similar directions, whereas smaller values indicate different directions.
Results.
We show the results in Fig. 7 (extended by standard deviations in Tab. 5 in the appendix). The results show very high similarity (0.98) between the update matrix and the $k=1$ approximation for GPT2-XL. For larger $k$ values the similarity drops significantly. For GPT-J, the similarity with $k=1$ is lower (0.77), but we have a moderate similarity (0.45) with $k=2$. Here too, the similarity values drop when $k>2$. For LLAMA3, the values are much lower (0.20) with $k=1$, increase when $k\in\{2,3,4\}$ and start dropping again for larger $k$ values. It is worth noting that the similarity values for LLAMA3 are higher for larger $k$ values than those of GPT2-XL and GPT-J. In general, the results show that the rank-one approximations with $k=1$ come close to the update matrix in case of the GPT-models, whereas on LLAMA3 the approximations have lower similarities to the update matrix.
Figure 7: The maximum cosine similarity values between vectors of the update matrix $W_{N}$ and the rank-one approximations $\tilde{W}^{(k)}_{V_{i}}$.
7.3 Reversal
The results from the previous section suggest that the editing information might be localized at the first few rank-one approximations of $W^{\prime}_{V}$, and that the original object before editing might still be encoded in bottom-rank approximations of $W^{\prime}_{V}$. This observation encourages us to investigate replacing the edited matrix $W^{\prime}_{V}$ by its bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$. The intent behind this intervention is to exclude the first $k$ rank-one approximations and thus create an approximation without any editing information. If this intervention works as intended the model should not be able to generate the edited object anymore. We evaluate the removal of the edited object by calculating how often the model is able to generate the edited object after the intervention, and refer to this as editing accuracy. Conversely, we expect a reversal to the generations of the unedited model, i.e, an increase in reversal accuracy.
Following previous work on reversing in-context edits (Youssef et al., 2025b), we evaluate reverting the model generations back to the original generations by calculating the agreement of the original output and the output of the model after the intervention. Editing and reversal accuracy are calculated as $\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\hat{y_{i}}=y_{i})$, where $y_{i}$ is the original output, and $\hat{y_{i}}$ are edited and reverse-edited output. Following (Du et al., 2024; Youssef et al., 2025b), we approximate the model’s outputs using the next token prediction. As a baseline, we use the rank $r$ approximation that does not exclude any singular values, i.e., we set $k=0$. Here, we use 310 instances, uniformly sampled from 31 relations.
Figure 8: The number of unique predictions with standard deviation when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ with $k\in\{0,\ldots,15\}$ with a set of 100 examples from wikitext-103 as inputs. Edited weights lead to more unique predictions. This finding can be used to identify edited weights.
Results.
The results in Tab. 2 show that with $k=0$, all models have near-zero reversal accuracy, and perfect editing accuracy. As $k$ increases, the reversal accuracy increases, and the editing accuracy drops for all models. Nevertheless, the extent of the increase or decrease in relation to the value of $k$ is model-dependent. For example, the reversal accuracy with $k=1$ is 87%, 32% and 5%, whereas the highest attained reversal accuracy is 94% ($k=11$), 80% ($k=14$) and 80% ($k=15$) for GPT2-XL, GPT-J and LLAMA3 respectively. We also notice that despite the high reversal accuracy, the editing accuracy never reaches 0%, especially on the larger GPT-J and LLAMA3 models. This indicates that some edits maybe more difficult to remove/reverse than others. In general the results show that the bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ can be used to remove the edit and retrieve the original outputs with high accuracy.
Qualitative analysis.
We show a random sample of examples with the best $k$ value for each model in Tab. 3. We generate 5 tokens given the input using greedy decoding. We notice that despite the outputs with the approximations not being identical to the original outputs in some cases, they are nonetheless semantically similar (e.g., soccer player/footballer, New York Mets/Jets, they/he). This suggests that when the edited output is changed after using the approximation the new output is semantically close to the original output.
Number of unique predictions.
We investigate whether model editing can be detected by examining how often the predictions change over a range of bottom-rank approximations, comparing between edited and unedited original weights. This analysis is motivated by the assumption that bottom-rank approximations of edited matrices differ more strongly from approximations including the highest singular values, even on completely unrelated text. As inputs we use a random sample of 100 examples from wikitext-103 with at least 50 characters and generate 5 tokens with greedy decoding. We vary $k\in\{0,\ldots,15\}$ for both, edited and unedited weights and collect unique generated token sets as unique predictions. The results in Fig. 8 show that bottom-rank approximations with edited weights lead to more unique predictions on average compared to unedited weights. For example, with GPT-J we have 1.37 predictions on average with unedited weights, but 2.46 predictions with edited weights. With LLAMA3 the gap is smaller (1.36 vs. 1.84). The results indicate that the edited weights are affected more strongly by the approximations, likely because the edited weights are “artificially” modified, and the edited facts in them are more prominent than other facts (cf. Sec. 6). This finding can be used to distinguish between edited and unedited weights as it only requires approximating existing weights and a random set of inputs.
8 Related Work
KEs can be categorized as either parameter-modifying, i.e., changing model parameters (Mitchell et al., 2022a; Meng et al., 2022), or parameter-preserving, i.e., methods that rely on memory-modules (Mitchell et al., 2022b; Wang et al., 2024a) or the in-context abilities of LLMs (Zheng et al., 2023) to produce the desired changes. Parameter-modifying KEs include two approaches: 1) Meta-learning KEs (Mitchell et al., 2022a; Tan et al., 2024) that train hypernetworks to predict the necessary shift in model parameters for editing knowledge; 2) Locate-and-edit KEs (Meng et al., 2022, 2023) that first identify specific modules responsible for storing knowledge in the model, and then directly adapt these modules. Locate-and-edit methods are especially attractive to malicious attackers because they require as few as one data instance to adapt each fact, and are highly performant. Our work focuses on ROME, a locate-and-edit KE, since recent work (Youssef et al., 2025a) shows that ROME is widely used in research on malicious knowledge editing.
Malicious knowledge editing.
Recent work (Youssef et al., 2025a) shows that knowledge editing methods can be used maliciously to implant backdoors (Li et al., 2024b), spread misinformation (Ju et al., 2024), bias (Chen et al., 2024), and jailbreak LLMs (Hazra et al., 2024). Particularly, Youssef et al. (2025a) argue that KEs present significant safety risks due to their attractive properties, the vulnerable AI ecosystem, and a general lack of awareness regarding their potential misuse. To date, limited work addresses countermeasures against model editing, with existing approaches primarily framing the problem as classification. These efforts focus on distinguishing between edited and unedited facts (Youssef et al., 2025c) and identifying different types of edits (Li et al., 2024a). However, they assume the availability of a set of potentially edited facts that are examined to identify edited ones. In this work, we address this limitation by first inspecting model weights for editing patterns, predicting the edited relation, and inferring the edited object based on model weights. Reversing edits has been limited to in-context edits (Youssef et al., 2025b). Our work complements this line of work by introducing an effective approach for reversing ROME-edits.
9 Conclusion
In this work, we studied the traces that editing with ROME, a prominent KE, leaves in LLMs, and investigated several approaches to trace and reverse ROME-edits. Our experiments spanned: (i) identifying edited weights, (ii) predicting the edited relation, (iii) inferring the edited object from model weights, and (iv) reversing edits using bottom-rank approximations. We showed how specific properties of ROME, i.e., the rank-one update and overfitting to the edited object, can be made use of to identify edited weights and gain more information about the edits. Our work contributes to transparent and safe AI by relaxing many assumptions from previous work on detecting knowledge, and paves the way for future research on additional methods.
Appendix A Limitations
Throughout our experiments, we assumed that only one edit is conducted with ROME, i.e., each edited matrix corresponded only to one edited fact. This corresponds to the setting in the original paper, and is not unrealistic, since a malicious attacker might be interested in keeping the edits minimal. Nonetheless, more than one edit can be conducted simultaneously and developing similar countermeasures to batch-editing methods would necessitate adapting the evaluation metrics to account for several edits.
Appendix B Autoregressive Transformers
A Transformer language model can be seen as a function $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ that maps an input $\mathbf{x}=(x_{1},...,x_{N})$ that consists of $N$ tokens to an output token $y\in\mathcal{Y}$. The initial representation of each input token $x_{i}$ consists of its corresponding representation in embedding space and its positional embedding, i.e., $h_{i}^{0}=encode(x_{i})+pos(x_{i})$ and $h_{i}^{0}\in\mathbb{R}^{d}$. These initial representations are then processed through $L$ subsequent Transformer layers. In each Transformer layer $l\in\{1,...,L\}$, the representations from the previous layers are processed using multi-head self-attention (MHSA) and MLP layers as follows:
$h_{i}^{l}=a_{i}^{l}+m_{i}^{l}+h_{i}^{l-1}$ (6)
$a_{i}^{l}=MHSA(h_{1}^{l-1},...,h_{i}^{l-1})$ (7)
$m_{i}^{l}=\sigma(W_{K}^{l}(a_{i}^{l}+h_{i}^{l-1}))W_{V}^{l}$ (8)
where $\sigma$ is a non-linear function, and $W_{K},W_{V}\in\mathbb{R}^{e\times d}$. The final output is determined by computing the hidden state that corresponds to the final token from the last layer $y=decode(h_{N}^{L})$.
Appendix C Additional Results
In this section, we provide more results. Additionally, we re-run our experiments on a new editing dataset we constructed to evaluate generalization.
Tab. 4 shows the relations used in our experiments. Tab. 5 shows the maximum cosine similarity values between vectors of the update matrix $W_{N}$ and the vectors of $\tilde{W}^{k}_{V_{i}}$ for different $k$ values.
Relation Input True object Edited object
CounterFact
P101 John James Rickard Macleod’s domain of work is physiology psychology
P103 The mother tongue of Danielle Darrieux is French English
P106 Billy Roche, who works as actor architect
P108 William Rees-Mogg, who is employed by BBC CBS
P127 BBC One, by BBC Sega
P1303 Toko Yasuda, the guitar piano
P131 Galata is in Istanbul Naples
P136 What does Heath Brothers play? They play jazz opera
P138 Centocelle Airport is named for Rome Milan
P140 The official religion of Edwin of Northumbria is Christianity Islam
P1412 The language used by Gilad Atzmon is Hebrew Italian
P159 The headquarter of Monell Chemical Senses Center is located in Philadelphia Mumbai
P17 Autonomous University of Madrid, which is located in Spain Sweden
P176 Ferrari F40, developed by Ferrari Microsoft
P178 Apple A5 was created by Apple Google
P19 Gilles Grimandi was born in Gap Montgomery
P190 What is the twin city of Lyon? It is Beirut Manila
P20 Charles Alfred Pillsbury expired at Minneapolis Berlin
P27 Mahmoud Fawzi has a citizenship from Egypt Germany
P276 Inner Circle railway line can be found in Melbourne Singapore
P30 Pidgeon Island belongs to the continent of Antarctica Asia
P364 The original language of The Icelandic Dream was Icelandic Tamil
P37 In Northwest Territories, an official language is English Tamil
P39 Robert William Muench is a bishop pope
P407 Mama Corsica was written in French Dutch
P413 Percy Snow, the linebacker goaltender
P449 The Loner was released on CBS HBO
P495 Shree Pundalik, created in India Sweden
P641 Andreas Ivanschitz professionally plays the sport soccer football
P740 Anaal Nathrakh, that was created in Birmingham Philadelphia
P937 Leonardo Balada found employment in Pittsburgh Paris
Yago
P112 The founder of Cabinn Hotels is Niels Fennet Toby Neugebauer
P131 The location of Nara Institute of Science and Technology is Japan Oran
P1405 The belief system of Al-Aziz Muhammad is Sunni Islam Anglicanism
P170 The artist of the painting The Marriage of the Virgin is Raphael Georges Braque
P171 The parent taxon of Puccinia recondita is Puccinia Microchiroptera
P178 The developer of Grand Theft Auto V is Rockstar London High Voltage Software
P200 The river Havel flows into Elbe Ōhura River
P238 The IATA code of Bankstown Airport is YSBK KGTB
P27 The nationality of Giulio Paradisi is Italy Hungary
P276 The location of the historical event Second Battle of Zurich is Zürich Constantinople
P463 The band of Freddie Mercury is Queen Love
P57 The director of Labyrinth of Flames is Katsuhiko Nishijima Carlo Vanzina
P840 The story of 24 is set in New York City Los Angeles
P915 The filming location of More Than Life at Stake is Poland France
P921 The subject of The Good Terrorist is terrorism social theory
Table 4: The relations we use in our experiments alongside examples.
Figure 9: Increase in row-wise cosine similarity of $W_{n}$ after editing. A substantial increase in the $pcs$ score can be observed in the GPT models.
Figure 10: Accuracy in generating the edited object based on the edited matrix when training different layers. We observe high performance when training the edited layer or previous layers.
Figure 11: The average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B
Max. Sim. Std Max. Sim. Std Max. Sim. Std
1 0.98 0.08 0.77 0.21 0.2 0.24
2 0.07 0.03 0.45 0.22 0.37 0.35
3 0.11 0.06 0.06 0.07 0.25 0.24
4 0.07 0.06 0.02 0.02 0.29 0.25
5 0.01 0.02 0.06 0.05 0.15 0.16
6 0.02 0.02 0.06 0.03 0.11 0.08
7 0.02 0.01 0.03 0.04 0.12 0.14
8 0.0 0.0 0.01 0.01 0.11 0.11
9 0.02 0.01 0.02 0.02 0.05 0.07
10 0.02 0.02 0.02 0.02 0.04 0.04
11 0.03 0.02 0.01 0.01 0.04 0.06
12 0.01 0.01 0.01 0.01 0.05 0.07
13 0.01 0.01 0.01 0.01 0.03 0.04
14 0.01 0.02 0.01 0.01 0.03 0.04
15 0.01 0.01 0.03 0.02 0.04 0.05
Table 5: The maximum cosine similarity values between vectors of the update matrix $W_{N}$ and the vectors of $\tilde{W}^{k}_{V_{i}}$ for different $k$ values.
Relation GPT2-XL GPT-J-6B META-LLAMA-3-8B
$pcs$ std $pcs$ std $pcs$ std
unedited 0.000091 NA 0.000194 NA $\approx$ 0.0 NA
P101 0.0067619231 0.0039638884 0.0018001686 0.0008643679 -3.756e-07 1.849e-07
P103 0.0059509007 0.00277431 0.0016752047 0.0006458259 -4.108e-07 1.845e-07
P106 0.0066805076 0.0026219752 0.0019167275 0.0007191846 -3.927e-07 2.394e-07
P108 0.0069100604 0.0031947018 0.0018263827 0.0007619029 -3.567e-07 1.895e-07
P127 0.0102751931 0.0051715499 0.0024409223 0.0010431761 -4.059e-07 1.763e-07
P1303 0.0076706613 0.0030963009 0.0020462978 0.0008670002 -3.889e-07 1.82e-07
P131 0.0098702735 0.0055993183 0.0045001531 0.0212773146 -3.732e-07 1.966e-07
P136 0.0071806164 0.0031261909 0.0019411065 0.0006443891 -4.322e-07 1.257e-07
P138 0.0137165404 0.0086577839 0.005331668 0.025452119 -4.461e-07 1.838e-07
P140 0.0078358574 0.0062726236 0.0020133093 0.0009927367 -3.933e-07 2.257e-07
P1412 0.006187233 0.0026656243 0.001784752 0.0005955094 -3.956e-07 1.998e-07
P159 0.008386647 0.0050185373 0.00218822 0.0009591461 -3.391e-07 2.753e-07
P17 0.009551276 0.0044502795 0.0026032751 0.0010471037 -3.552e-07 2.959e-07
P176 0.0108912224 0.0055227291 0.0019816675 0.00099757 -4.325e-07 1.372e-07
P178 0.0117567854 0.0110017566 0.0021579888 0.001029392 -4.225e-07 1.659e-07
P19 0.0074281091 0.0030904648 0.0021313282 0.0009460897 -3.971e-07 1.827e-07
P190 0.0178613561 0.0165129782 0.0029858089 0.0012192149 -4.938e-07 1.712e-07
P20 0.007044959 0.0032487115 0.001820856 0.0009291652 -4.047e-07 1.481e-07
P27 0.0072250371 0.0033331825 0.0018882287 0.0006491209 -3.945e-07 1.818e-07
P276 0.0096739383 0.00348082 0.0021771877 0.000703454 -3.604e-07 2.357e-07
P30 0.0104348788 0.0078027795 0.0028299001 0.0011365804 -4.144e-07 2.235e-07
P364 0.0055471736 0.0028574185 0.0016467369 0.0006391143 -4.365e-07 2.311e-07
P37 0.0073569546 0.0043815362 0.001865609 0.0008194575 -4.034e-07 2.107e-07
P39 0.0074461334 0.0031980671 0.0019889568 0.0008512599 -3.884e-07 2.081e-07
P407 0.0067631754 0.0027889247 0.0018608728 0.0007570319 -3.982e-07 1.961e-07
P413 0.0069200734 0.0031209039 0.0019079371 0.0007796126 -3.95e-07 1.713e-07
P449 0.0071908987 0.0040715897 0.0020522842 0.0009013923 -3.81e-07 1.799e-07
P495 0.0088219754 0.0069427193 0.0022077068 0.0009163789 -3.528e-07 2.456e-07
P641 0.0056973715 0.0026820171 0.0017082412 0.0007183695 -3.228e-07 2.105e-07
P740 0.0086965727 0.0042147134 0.0040305918 0.0150149984 -3.625e-07 2.245e-07
P937 0.0069481413 0.0047418696 0.0018729484 0.0008598884 -3.912e-07 1.558e-07
Table 6: Pair-wise cosine similarity ($pcs$) scores with different relations from CounterFact.
C.1 Yago Dataset
Dataset construction.
We use the knowledge base YAGO 4.5 Suchanek et al. (2024) to create an editing dataset that contains more diverse relations than CounterFact. YAGO 4.5 merges the taxonomy of Wikidata with the taxonomy of Schema.org to create a consistent knowledge base. We manually selected a set of diverse relations from YAGO and extracted the corresponding subject and object pairs. We filtered out relations with less than 1000 subject/object pairs. Afterwards we used DeepSeek R1 DeepSeek-AI et al. (2025) to generate editing prompts and paraphrased prompts for each relation. We manually checked the correctness of the generated prompts. YAGO 4.5 is licensed under a Creative Commons Attribution 4.0 International License (CC BY 4.0), and we will publish the dataset under the same license. We use 15 relations (shown in Tab. 4) from the newly constructed dataset in our experiments.
Results on Yago.
Fig. 13 shows the direction distribution of row vectors in the update matrix. Fig. 14 shows the average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers. Fig. 15 shows the increase in row-wise cosine similarity after editing. Tab. 7 shows the results for predicting the edited relation. Fig. 12 shows the accuracy of inferring the edited objects based on the edited weights.
GPT2-XL GPT-J META-LLAMA-3-8B
#Classes Baseline Accuracy Std Accuracy Std Accuracy Std
2 50.6 98.4 1.95 97.0 3.08 95.0 2.83
3 30.53 99.87 0.3 97.87 1.45 94.27 3.35
5 19.6 97.52 1.04 95.68 2.6 90.08 3.77
10 10.32 94.76 0.96 88.64 2.35 82.24 5.2
15 6.77 92.32 1.05 83.87 1.44 74.27 1.44
Table 7: Accuracy for predicting the edited relation based on low-dimensional representations of the edited matrices using a logistic regression classifier. We experiment with different number of relations (#Classes). The relations used are from the Yago dataset.
$k$ GPT2-XL GPT-J-6B META-LLAMA-3-8B
Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$ Reversal Acc. $\uparrow$ Editing Acc. $\downarrow$
0 0.00 $\pm$ 0.00 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 95.71 $\pm$ 20.40 1.43 $\pm$ 11.95 91.43 $\pm$ 28.20
1 92.86 $\pm$ 25.94 5.71 $\pm$ 23.38 24.29 $\pm$ 43.19 61.43 $\pm$ 49.03 5.71 $\pm$ 23.38 82.86 $\pm$ 37.96
2 95.71 $\pm$ 20.40 2.86 $\pm$ 16.78 68.57 $\pm$ 46.76 10.00 $\pm$ 30.22 28.57 $\pm$ 45.50 41.43 $\pm$ 49.62
3 94.29 $\pm$ 23.38 1.43 $\pm$ 11.95 74.29 $\pm$ 44.02 8.57 $\pm$ 28.20 37.14 $\pm$ 48.67 34.29 $\pm$ 47.81
4 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 74.29 $\pm$ 44.02 8.57 $\pm$ 28.20 50.00 $\pm$ 50.36 20.00 $\pm$ 40.29
5 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 4.29 $\pm$ 20.40 54.29 $\pm$ 50.18 17.14 $\pm$ 37.96
6 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 84.29 $\pm$ 36.66 4.29 $\pm$ 20.40 58.57 $\pm$ 49.62 17.14 $\pm$ 37.96
7 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 82.86 $\pm$ 37.96 0.00 $\pm$ 0.00 65.71 $\pm$ 47.81 10.00 $\pm$ 30.22
8 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 84.29 $\pm$ 36.66 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 8.57 $\pm$ 28.20
9 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 7.14 $\pm$ 25.94
10 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 80.00 $\pm$ 40.29 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
11 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 80.00 $\pm$ 40.29 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
12 94.29 $\pm$ 23.38 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 0.00 $\pm$ 0.00 70.00 $\pm$ 46.16 8.57 $\pm$ 28.20
13 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 0.00 $\pm$ 0.00 72.86 $\pm$ 44.79 7.14 $\pm$ 25.94
14 97.14 $\pm$ 16.78 0.00 $\pm$ 0.00 78.57 $\pm$ 41.33 1.43 $\pm$ 11.95 70.00 $\pm$ 46.16 7.14 $\pm$ 25.94
15 95.71 $\pm$ 20.40 0.00 $\pm$ 0.00 81.43 $\pm$ 39.17 1.43 $\pm$ 11.95 71.43 $\pm$ 45.50 5.71 $\pm$ 23.38
Table 8: Reversal/Editing accuracy on Yago with different $r-k$ approximations of $W^{\prime}_{V}$. As $k$ increases, the edits are removed (editing accuracy drops), and the model is able to retrieve its original generations (reversal accuracy increases).
Input $k$ Edited Object Orig. Output App. Output
GPT2-XL
The founder of Cabinn Hotels is 11 Toby Neugebauer a former U.S a man who has been
The river Ergolz flows into 11 Jiu River the Black Sea. The the Black Sea. The
The band of Jeff Tweedy is 11 Anthrax a band of friends. a band of friends.
The band of Iain Matthews is 11 Dire Straits a band of Iain a band of the future
The river Havel flows into 11 Ōhura River the Danube, and the Danube, and
GPT-J
The founder of Tune Hotels is 6 Luís I of Portugal a man who has been a man who has been
The director of The Mirror is 6 Polly Draper a man who has been a man who has been
The director of Darkman is 6 Claudio Fragasso a man who has been a man who has been
The story of 24 is set in 6 Los Angeles the year 2024, and the year 2401,
The subject of Goryeosa is 6 orphan the life of the Buddha the story of the life
META-LLAMA-3-8B
The artist of the painting Allegory of Vices is 11 Mary Cassatt unknown. The painting was Mary Cassatt Mary
The artist of the painting Religious Procession in Kursk Province is 11 Giulio Romano Ivan Ivanovich Shish Ivan Ivanovich Shish
The river Melbbach flows into 11 Inn the river Main in the the river Inn at the
The subject of Net Voyne! is 11 international relations the Internet and its impact the Internet and its impact
The director of Brave Command Dagwon: The Boy with Crystal Eyes is 11 Sally Potter back with a new anime a 1996 anime
Table 9: Model outputs when using bottom-rank approximations $\tilde{W^{\prime}}_{V}^{(r,k)}$ on a random set of facts. We use the best $k$ for each model. The examples show that the model outputs with approximations (Approx. Output) are semantically close to the original/unedited outputs (Orig. Output).
Figure 12: Accuracy in generating the edited object based on the edited matrix when training different layers. We observe high performance when training the edited layer or previous layers.
Figure 13: Percentage of row vectors in the update matrix $W_{N}$ having the same direction or opposite directions. More than 80% of the vectors have the same direction in the GPT models. Yago Dataset.
Figure 14: The average pairwise cosine similarity ($pcs$) of edited and unedited matrices from different layers (Yago Dataset).
Figure 15: Increase in row-wise cosine similarity of $W_{n}$ after editing with the Yago dataset. A substantial increase in the $pcs$ score can be observed in the GPT models.
Appendix D Implementation Details
Tab. 10 shows the dimensionality of the edited matrices in each model. Sec. D.1 describes the compute resources.
Model Edited Matrix Dim.
GPT2-XL $6400\times 1600$
GPT-J-6B $16384\times 4096$
META-LLAMA-3-8B $14336\times 4096$
Table 10: The dimensionalities of the edited matrices for different models.
Dataset License
CounterFact Meng et al. (2022) MIT License
YAGO Suchanek et al. (2024) CC BY 4.0
Table 11: The datasets we use in this work and their licenses.
D.1 Compute Resources
We ran all of our experiments were on an NVIDIA A100 GPU with 80GB. We started our experiments by editing the 3 models and storing the weight matrices. Editing took roughly 7 GPU days. The experiments for predicting the edits took roughly 2 GPU days. The experiments for inferring the edits from edited weights took roughly 7 GPU days. The experiments for reversing edits took roughly 5 GPU days.